# Dark energy as a geometrical effect in geodetic brane gravity

### v1

@@ Abstract
Within the framework of the modified geodetic brane gravity, conformed by the Regge-Teitelboim model and enhanced with a linear term in the extrinsic curvature of the brane, the possibility that under an FRW geometry this theory emulates the so-called dark energy is discussed. The cosmological behavior of this model displays a self-(non-self)-accelerated expansion of this universe which is caused by a combination of usual matter and gravitational geometric effects controlled by a $\beta$ parameter that accompanies the correction $K$ term. Indeed, the self-accelerated branch, provided by the trace $K$ model raises the question of whether the extrinsic curvature correction terms might be suitable for dark energy candidates. We discuss the analytical expression obtained for $\rho_{\text{\tiny dark}}$ in addition to the main cosmological parameters such as the state parameter $\omega_{\text{\tiny eff}}$ and the deceleration parameter $q$. Moreover, when we call for the contribution of dark radiation-like energy to be switched off, $\Omega_{dr,0}\to 0$, we find the same acceleration behavior, as well as the same dark energy content provided by the DGP theory.
@@ I Introduction
Any attempt to solve the still unexplained mystery of the dark matter/energy, as long as it is based on fundamental geometric symmetries which play a central role in respected gravitational theories, is welcome. The dark energy nature constitutes a tantalizing question to which the cosmology and the particle physics are attempting to provide a compelling answer. Following this line of reasoning, within the geodetic brane gravity based in the Regge-Teitelboim (RT) model, we would like to explore a possible explanation to such a puzzle where extrinsic curvature terms can provide findings that reproduce many of the characteristics that exhibits the so-called dark matter/energy. The RT gravity theory generalizes General Relativity (GR) in a particular theoretical direction [1]. The theory is based on a simple particle/string-inspired assumption that our Universe is a 4-dimensional surface isometrically embedded, and evolving, in a higher-dimensional flat ambient spacetime. In this view, our 4-dimensional spacetime is regarded as the trajectory (worldvolume) of a 3-dimensional extended object, probing a higher-dimensional flat ambient space. Within this assumption, one must not overlook the local isometric embedding theorem [2, 3]. The RT action is identical to the Einstein-Hilbert (EH) action of GR but, with the crucial difference that the embedding functions $X^{\mu}(x^{a})$ rather than the metric $g_{ab}$ defined of the worldvolume, are elevated to the level of field variables. The field equations result in the Einstein ones but contracted with the extrinsic curvature tensor of the embedding which represents a generalization of the Einstein equations in the sense that the RT equations are fulfilled for every solution of the Einstein equations. The theory has attractive properties such as the already mentioned built-in Einstein limit and in particular, within the braneworld context, a dark companion to any primitive energy density. Under the names of geodetic brane gravity or embedding gravity, the RT model has been widely investigated by many authors as an interesting formulation to describe gravity, cosmological implications in extra dimensions as well as a suitable alternative for quantization of gravity [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
@@
In this paper, within the extended objects framework, we attempt to go one step beyond pure RT gravity and reignite the discussion regarding the dark energy companion of the primitive $\rho$ for the RT cosmology as discussed in [22, 23, 24], by exercising the geometrical features contained in the trace $K$ model. Specifically, in this regard, we show that a kind of dark energy is a geometrical effect. Certainly, some braneworld models at low energies have become interesting scenarios providing late-time acceleration behaviors without the need to consider dark energy by adding in the extra mysterious matter. We observe that $\beta$, the parameter accompanying the trace $K$ model, is the only input parameter in our development that produces this effect. In this sense, with the claim of exploring the source of the dark energy, within the modified gravity theories context, one of the requirements is the built-in Einstein limit. On the basis of geodetic brane gravity, it is interesting and motivating to improve the theory by incorporating properly pathologies-free geometric terms so that physical effects from other more elaborated theories can be reproduced. Despite the bulk being flat Minkowski, there are local effects on the brane via conserved quantities associated with the Poincaré invariance, and the correction $K$ term which can be thought of as a remnant of the existence of a 5-dimensional EH action. We find that our proposal, when applied to a homogeneous and isotropic universe that has an FRW geometry, leads to a suggestive modification of the Friedmann equation which includes an additional term, $\rho_{\text{\tiny dark}}$, in the effective energy density and that it is treated in the manner that the primitive $\rho$. The presence of such $\rho_{\text{\tiny dark}}$ originates mainly from the extrinsic curvature correction term $K$ and the matter confined to the braneworld. The trace $K$ model requires careful observation; needles to say such a term resembles the Gibbons-Hawking-York (GHY) counterterm but with the essential difference that in our approach $K$ contains time derivatives of the field variables that contribute to the dynamics of the brane. It is worth mentioning this model has cropped up in other contexts ranging from lipid membranes to particle physics and brane-world gravity [25, 26, 27].
@@
In the first part of our work, our strategy is, under an FRW geometry on the brane, to make use of the invariance under reparametrizations of the model and identify conserved quantities that help to integrate the equation of motion, and thus lead us to identify a Friedmann type equation. By invoking, and imposing, the traditional point of view of the cosmology for the Friedmann equation, we may identify an effective energy density where $\rho_{\text{\tiny dark}}$ is a companion piece of the primitive energy density $\rho$. Indeed, we can express the dark energy contribution in terms of the associated extrinsic curvature correction parameter, $\beta$, and the confined matter, producing a self-(non-self)-acceleration behavior of this type of universe. One could wonder if the answer to the tantalizing question of dark energy is related to the inherent extrinsic curvature and its embedding properties, as was perceived in [28], so our approach suggests that is the answer. Subsequently, we analytically compute important cosmological parameters such as the equation of state parameter and the deceleration parameter. These confirm the existence of a type of dark energy density and the existence of a self-(non-self)-acceleration behavior through the emergence of Big Chill and Big Bounce histories of this universe. It is worth mentioning that this model suffers from a fine-tuning problem, so one must rely on numerical analysis.
@@ III Dark energy from modified geodetic brane cosmology
with $\omega$ being a constant, the fingerprint of the background spacetime. In fact, $\omega$ plays a double duty. On the one hand, it is related to the conserved bulk energy and, on the other hand, this parameterizes the deviation from the Einstein limit when $\omega=0$ in addition to $\beta=0$. In fact, after some algebra and solving an involved cubic equation, (20) leads to an effective Friedmann equation [19].
@@
We remark in passing that, the non-standard Friedmann equation in our approach is provided by $\dot{a}^{2}+k=f(a;\Omega_{k,0},\Omega_{m,0},\Omega_{\beta,0},\Omega_{dr,0})a^{2}$ where the energy density parameters are given in (28), and we have additionally introduced the energy density parameter $\Omega_{k,0}:=-\frac{k}{H_{0}^{2}}$[19]. One can go one step further and cast this Friedmann-like equation in the suggestive form $\dot{a}^{2}+U(a;\Omega_{s})=0$ where $U$ stands for an effective potential parameterized by the density energies $\Omega_{s}$, (28), [19]. This fashion is convenient since it allows the system to be faced as an equivalent one-dimensional non-relativistic mechanical problem in which the total mechanical energy is zero. Certainly, we have
@@
$-\left.\left.\frac{2\omega_{m,\beta,z}\left[\left(4(1+z)\Omega_{dr,0}+\Omega_{m,0}\Omega_{\beta,0}\right)\omega_{m,\beta,z}^{2}+\Omega_{m,0}\Omega_{m,\beta,dr,z}\right]\sin(\varphi_{m,\beta,dr,z})}{\sqrt{\omega_{m,\beta,z}^{6}-\Omega_{m,\beta,dr,z}^{2}}}\right]\right\}$
@@ IV Concluding remarks
In this paper, we have analyzed the dark energy content within the geodetic brane gravity context enhanced with an extrinsic curvature term associated with the shape of the brane. We provide evidence that such dark energy can originate from the correction of extrinsic curvature terms appearing in the action functional that models our Universe. To show this we have considered a FRW geometry defined onto our braneworld probing a $5$-dimensional Minkowski background spacetime. Our approach is free of a cosmological constant with the idea of highlighting the role that the correction term $K$ plays in development. In this sense, it plays the role of a cosmological constant providing self-(and non-self) acceleration behavior. One can think of the $K$ term as a relic or a possible fingerprint of the existence of a $5$-dimensional EH action that disappears at some time. By assuming a traditional point of view for the Friedmann equation, we observed that the regular matter is accompanied by an atypical piece, termed $\rho_{\text{\tiny dark}}$. In our development, the intriguing dark energy effect results as a geometrical achievement originated by the model under consideration. What is notable is to observe how the correction terms associated with the extrinsic curvature of the brane can reproduce many of the general characteristics of the so-called dark energy.
Our approach is not intended to be realistic, but on the contrary, it aims to be the promoter of extrinsic curvature corrections that help to understand several physical phenomena ranging from condensed soft matter to cosmology and particle physics. Because of this, perhaps the inclusion of some matter fields, such as the cosmological constant, can take on a more realistic meaning. A consequent extension of our work presented here is the incorporation of other attractive extrinsic curvature terms such as those provided by the so-called Lovelock brane gravity theory. For instance, the Gibbons-Hawking-York-Mayers type term, $K^{3}-2KK_{ab}K^{ab}+3K^{a}{}_{b}K^{b}{}_{c}K^{c}{}_{a}$, [30]. We believe that such an addition could improve or enrich the cosmological findings in this braneworld context. This will be reported elsewhere.
@@ Acknowledgements.
The authors are grateful to R. Cordero, J. A. Méndez Zavaleta and M. Cruz for discussions and helpful suggestions. ER acknowledges encouragement from ProDeP-México, CA-UV-320: Álgebra, Geometría y Gravitación. JCNZ acknowledges support from a CONAHCYT-México graduate fellowship. GC acknowledges support from a Postdoctoral Fellowship by Estancias Posdoctorales por México 2023(1)-CONAHCYT . Also, ER thanks partial support from Sistema Nacional de Investigadores, México.

### v2

@@ Abstract
Within the framework of the modified geodetic brane gravity, conformed by the Regge-Teitelboim model and enhanced with a linear term in the extrinsic curvature of the brane, the possibility that under an FRW geometry this theory emulates the so-called dark energy is discussed. The cosmological behavior of this model displays a self-(non-self)-accelerated expansion of this universe which is caused by a combination of usual matter and gravitational geometric effects controlled by a $\beta$ parameter that accompanies the correction $K$ term. Indeed, the self-accelerated branch, provided by the trace $K$ model raises the question of whether the extrinsic curvature correction terms might be suitable for dark energy candidates. We discuss the analytical expression obtained for $\rho_{\text{\tiny dark}}$ in addition to the main cosmological parameters such as the state parameter $\omega_{\text{\tiny eff}}$ and the deceleration parameter $q$. Moreover, when we call for the contribution of dark radiation-like energy to be switched off, $\Omega_{dr,0}\to 0$, we find the same acceleration behavior, as well as the same dark energy content provided by the DGP theory. The relationship of our findings to the analysis for $\rho_{\text{\tiny dark}}$ performed by Davidson and Gurwich within the unified brane cosmology is briefly discussed.
@@ I Introduction
Any attempt to solve the still unexplained mystery of the dark matter/energy, as long as it is based on fundamental geometric symmetries which play a central role in respected gravitational theories, is welcome. The dark energy nature constitutes a tantalizing question to which the cosmology and the particle physics are attempting to provide a compelling answer. Following this line of reasoning, within the geodetic brane gravity based in the Regge-Teitelboim (RT) model, we would like to explore a possible explanation to such a puzzle where extrinsic curvature terms can provide findings that reproduce many of the characteristics that exhibits the so-called dark matter/energy. The RT gravity theory generalizes General Relativity (GR) in a particular theoretical direction [1]. The theory is based on a simple particle/string-inspired assumption that our Universe is a 4-dimensional surface isometrically embedded, and evolving, in a higher-dimensional flat ambient spacetime. In this view, our 4-dimensional spacetime is regarded as the trajectory (worldvolume) of a 3-dimensional extended object, probing a higher-dimensional flat ambient space. Within this assumption, one must not overlook the local isometric embedding theorem [2, 3, 4]; indeed, in order to ensure the local existence of an embedding framework, at most $N=n(n+1)/2$ background flat dimensions are required to locally embed a $n$-metric. Clearly, for a $4$-dimensional brane, at most a $10$-dimensional flat ambient spacetime is needed, a number that can be reduced when the brane metric admits some Killing vector fields as fortunately occurs in the case of an FRW brane geometry. In this sense, as one learns for example in [3], with $N=5$ background flat dimensions it is sufficient. The RT action is identical to the Einstein-Hilbert (EH) action of GR but, with the crucial difference that the embedding functions $X^{\mu}(x^{a})$ rather than the metric $g_{ab}$ defined of the worldvolume, are elevated to the level of field variables. The field equations result in the Einstein ones but contracted with the extrinsic curvature tensor of the embedding which represents a generalization of the Einstein equations in the sense that the RT equations are fulfilled for every solution of the Einstein equations. The theory has attractive properties such as the already mentioned built-in Einstein limit and in particular, within the braneworld context, a dark companion to any primitive energy density. Under the names of geodetic brane gravity or embedding gravity, the RT model has been widely investigated by many authors as an interesting formulation to describe gravity, cosmological implications in extra dimensions as well as a suitable alternative for quantization of gravity [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
In this paper, within the extended objects framework, we attempt to go one step beyond pure RT gravity and reignite the discussion regarding the dark energy companion of the primitive $\rho$ for the RT cosmology as discussed in [23, 24, 25], by exercising the geometrical features contained in the trace $K$ model. Specifically, in this regard, we show that a kind of dark energy is a geometrical effect. Certainly, some braneworld models at low energies have become interesting scenarios providing late-time acceleration behaviors without the need to consider dark energy by adding in the extra mysterious matter. We observe that $\beta$, the parameter accompanying the trace $K$ model, is the only input parameter in our development that produces this effect. In this sense, with the claim of exploring the source of the dark energy, within the modified gravity theories context, one of the requirements is the built-in Einstein limit. On the basis of geodetic brane gravity, it is interesting and motivating to improve the theory by incorporating properly pathologies-free geometric terms so that physical effects from other more elaborated theories can be reproduced. A pioneer work in this direction has been performed in [26] within the framework of unified brane cosmology where Davidson and Co., find a dark energy expression representing a dark companion of the observable $\rho$. Despite the bulk being flat Minkowski, there are local effects on the brane via conserved quantities associated with the Poincaré invariance, and the correction $K$ term which can be thought of as a remnant of the existence of a 5-dimensional EH action. We find that our proposal, when applied to a homogeneous and isotropic universe that has an FRW geometry, leads to a suggestive modification of the Friedmann equation which includes an additional term, $\rho_{\text{\tiny dark}}$, in the effective energy density and that it is treated in the manner that the primitive $\rho$. The presence of such $\rho_{\text{\tiny dark}}$ originates mainly from the extrinsic curvature correction term $K$ and the matter confined to the braneworld. The trace $K$ model requires careful observation; needles to say such a term resembles the Gibbons-Hawking-York (GHY) counterterm but with the essential difference that in our approach $K$ contains time derivatives of the field variables that contribute to the dynamics of the brane. It is worth mentioning this model has cropped up in other contexts ranging from lipid membranes to particle physics and brane-world gravity [27, 28, 29].
@@
In the first part of our work, our strategy is, under an FRW geometry on the brane, to make use of the invariance under reparametrizations of the model and identify conserved quantities that help to integrate the equation of motion, and thus lead us to identify a Friedmann type equation. By invoking, and imposing, the traditional point of view of the cosmology for the Friedmann equation, we may identify an effective energy density where $\rho_{\text{\tiny dark}}$ is a companion piece of the primitive energy density $\rho$. Indeed, we can express the dark energy contribution in terms of the associated extrinsic curvature correction parameter, $\beta$, and the confined matter, producing a self-(non-self)-acceleration behavior of this type of universe. One could wonder if the answer to the tantalizing question of dark energy is related to the inherent extrinsic curvature and its embedding properties, as was perceived in [30], so our approach suggests that is the answer. On the other hand, the intricate analytical expression that we found to know $\rho_{\text{\tiny dark}}$, Eq. (20), coincides with that previously obtained by Davidson and Gurwich in [26] making use of a Dirac’s brane variation prescription within the unified brane theory framework. Subsequently, we analytically compute important cosmological parameters such as the equation of state parameter and the deceleration parameter. These confirm the existence of a type of dark energy density and the existence of a self-(non-self)-acceleration behavior through the emergence of Big Chill and Big Bounce histories of this universe. It is worth mentioning that this model suffers from a fine-tuning problem, so one must rely on numerical analysis.
@@ III Dark energy from modified geodetic brane cosmology
with $\omega$ being a constant. In fact, $\omega$ plays a double duty. On the technical side, it is related to the conserved bulk energy conjugate to the embedding time coordinate $t(\tau)$ and, on the physical side, within the unified brane cosmology [26], it parameterizes the deviation from the Randall-Sundrum brane cosmology and from general relativity when $\omega=0$ in addition to $\beta=0$. With an eye to cosmology, this last remark represents the main role of the $\omega$ constant. An additional remark is in order; as discussed in [26] $\omega$ is not necessarily a small quantity and its size will be fixed once the dark matter interpretation is fully established. Further, after some algebra and solving an involved cubic equation, (20) leads to an effective Friedmann equation [20].
@@
It is worth noting that this universe, free of a cosmological constant, does not provide relevant information at early times making it unrealistic at this stage. To show this it is convenient to explore the $\rho=0$ scenario which is dominated only by the parameters $\omega$ and $\beta$. By specializing the cubic equation (23) with $\rho=0$ and solving for $\mathcal{Z}$ in addition to focusing on the real solution, we are able to determine, after an expansion around $a=0$, that the significant contribution to $\rho_{\text{\tiny dark}}$ is provided by negative values of $\omega$, given by
$\rho_{\text{\tiny dark}}\simeq\frac{|\omega|^{2/3}}{a^{8/3}}.$ (32)
Therefore, the possibility of obtaining a premature model is not possible since this result, which agrees with the one discussed in [24] under similar conditions, this type of universe evolves similarly as a matter dominated FRW universe which does not represent a realistic universe. In this sense, it seems that the model only has an impact for late times since it reproduces an accelerated behaviour. We remark in passing that, the non-standard Friedmann equation in our approach is provided by $\dot{a}^{2}+k=f(a;\Omega_{k,0},\Omega_{m,0},\Omega_{\beta,0},\Omega_{dr,0})a^{2}$ where the energy density parameters are given in (28), and we have additionally introduced the energy density parameter $\Omega_{k,0}:=-\frac{k}{H_{0}^{2}}$[20]. One can go one step further and cast this Friedmann-like equation in the suggestive form $\dot{a}^{2}+U(a;\Omega_{s})=0$ where $U$ stands for an effective potential parameterized by the density energies $\Omega_{s}$, (28), [20]. This fashion is convenient since it allows the system to be faced as an equivalent one-dimensional non-relativistic mechanical problem in which the total mechanical energy is zero. Certainly, we have
@@
$-\left.\left.\frac{2\omega_{m,\beta,z}\left[\left(4(1+z)\Omega_{dr,0}+\Omega_{m,0}\Omega_{\beta,0}\right)\omega_{m,\beta,z}^{2}+\Omega_{m,0}\Omega_{m,\beta,dr,z}\right]\sin(\varphi_{m,\beta,dr,z})}{\sqrt{\omega_{m,\beta,z}^{6}-\Omega_{m,\beta,dr,z}^{2}}}\right]\right\},$
@@ IV Concluding remarks
In this paper, we have analyzed the dark energy content within the geodetic brane gravity context enhanced with an extrinsic curvature term associated with the shape of the brane. We provide evidence that such dark energy can originate from the correction of extrinsic curvature terms appearing in the action functional that models our Universe. To show this we have considered a FRW geometry defined onto our braneworld probing a $5$-dimensional Minkowski background spacetime. Our approach is free of a cosmological constant with the idea of highlighting the role that the correction term $K$ plays in development. In this sense, it plays the role of a cosmological constant providing self-(and non-self) acceleration behavior at late times. One can think of the $K$ term as a relic or a possible fingerprint of the existence of a $5$-dimensional EH action that disappears at some time. Regarding this point, on geometrical grounds, it is worth noting that including the $K$-term is only possible for timelike oriented hypersurfaces floating in a background spacetime. Indeed, we can not lift this restriction and consider an arbitrary co-dimension because it would be impossible to maintain the symmetry of the action functional. By assuming a traditional point of view for the Friedmann equation, we observed that the regular matter is accompanied by an atypical piece, termed $\rho_{\text{\tiny dark}}$. In our development, the intriguing dark energy effect results as a geometrical achievement originated by the model under consideration. What is notable is to observe how the correction terms associated with the extrinsic curvature of the brane can reproduce many of the general characteristics of the so-called dark energy.
Our approach is not intended to be realistic, but on the contrary, it aims to be the promoter of extrinsic curvature corrections that help to understand several physical phenomena ranging from condensed soft matter to cosmology and particle physics. Because of this, perhaps the inclusion of some matter fields, such as the cosmological constant, can take on a more realistic meaning. It would be interesting to extend the analysis of our cosmological findings and verify to what extent the results obtained are close to current observational data such as those provided by the $\Lambda$CDM model. This involves, for example, graphing the different $\Omega$s as functions of Log$[a]$ for comparison, but this would involve modifying some starting expressions. This is however a more deep numerical analysis that is beyond the scope of this note. A consequent extension of our work presented here is the incorporation of other attractive extrinsic curvature terms such as those provided by the so-called Lovelock brane gravity theory. For instance, the Gibbons-Hawking-York-Mayers type term, $K^{3}-2KK_{ab}K^{ab}+3K^{a}{}_{b}K^{b}{}_{c}K^{c}{}_{a}$, [32]. We believe that such an addition could improve or enrich the cosmological findings in this braneworld context. This will be reported elsewhere.
@@ Acknowledgements.
We would like to thank anonymous referees for constructive comments and for drawing our attention to some missing references. The authors are grateful to R. Cordero, J. A. Méndez Zavaleta and M. Cruz for discussions and helpful suggestions. ER acknowledges encouragement from ProDeP-México, CA-UV-320: Álgebra, Geometría y Gravitación. JCNZ acknowledges support from a CONAHCYT-México graduate fellowship. GC acknowledges support from a Postdoctoral Fellowship by Estancias Posdoctorales por México 2023(1)-CONAHCYT . Also, ER thanks partial support from Sistema Nacional de Investigadores, México.